what mass of sodium acetate, NaCH3COO, do we need to add to a 500 mL solution that contains 1.2g of acetic acid, CH3COOH, to make a buffered solution with pH 5.0. No value change when add sodium acrtate. Ka of acid is 1.76 x10-5
Answer #1

mixture of CH3COOH and CH3COONa act as acidic buffer

for such buffers

pH = pKa + log [CH3COONa] / [CH3COOH]

pH = 5.0

pka = - log Ka = - log 1.76 x 10-5 = 4.75

[CH3COOH] = (1.2 / 60.05) (1000 / 500) = 0.040 M

[CH3COONa] = ?

5.0 = 4.75 + log [COONa] / [0.040]

0.25 = log [CH3COONa] + 1.40

log [CH3COONa] = -1.15

[CH3COONa] = 0.071 M.

0.071 = (W / 82.03) (1000 / 500)

0.0355 = W / 82.03

W = 2.91 g

mass of CH3COONa = 2.91 g